# A new general quantum state verification protocol by the classical shadow method

### v3

@@ IV Directly partial shadow overlap protocol
Figure 3: Illustration of the averaged spectral gaps of $\nu(L)$ and $\nu(\Omega)$ for the SOP and DPSO protocol. The target states are Haar-random pure states, and the averaged spectral gaps are computed over $1000$ samples. The first figure shows that the averaged spectral gaps decreases as the number of qubits increases, and that averaged $\nu(\Omega)$ is consistently larger than the averaged $\nu(L)$. The second figure illustrates the averaged spectral gaps increase with the protocol level, and the averaged spectral gap $\nu(\Omega)$ exceeds the averaged $\nu(L)$ across all levels.
We have discussed in detail the differences between the PLM framework and the SOP. In the SOP, only Pauli-Z measurements are applied to $n-r$ qubits, which prevents the direct application of the SOP to certain states, such as GHZ states. Additionally, the operator $L$ appears somewhat unnatural and complex. To address these limitations, we aim to generalize the protocol by incorporating other types of Pauli measurements and selecting a more natural operator than $L$. Inspired by measurements used in [23], we propose a new protocol, called as the directly partial shadow overlap (DPSO) protocol, which follows the same core principles as the SOP and utilizes the classical shadow method.
@@ IV.1 Description of the verification procedure
In general, our DPSO protocol is similar with the SOP in [1]. As outlined in Protocol 3, we need to run $N$ times of Algorithm 2 to compute $N$ shadow overlaps $\hat{\omega}_{a}$. In Algorithm 2, the steps $1,2,3$ correspond to the first stage of the SOP, while the steps $4,5$ correspond to the second stage of classical post-processing. The procedure of Algorithm 2 is also illustrated in Figure 4.
@@
  • For the randomly selected $n-r$ qubits, we randomly apply Pauli-$X$, $Y$, or $Z$ measurements, rather than only the Pauli-$Z$ measurements.
@@
The last two differences are more important. The second modification enables us to handle a wider range of states with special structures, such as the GHZ states and stabilizer states. The third adjustment opens the possibility of constructing the optimal protocol, where the strategy operator achieves the optimal spectral gap. Hence, when the level $r=1$ and only Pauli-$Z$ measurements are applied, the DPSO protocol reduces to the SOP and is similar with the protocol in [23], then the DPSO protocol can also verify almost all Haar random pure states, just as the SOP [1]. Finally, it is important to emphasize that the measurement applied in each run of the DPSO protocol is not an LOCC measurement, but rather a randomized Pauli measurement. Therefore, the DPSO protocol should not be regarded as an adaptive version of the SOP.
@@
1:Label $n$ qubits of state $\rho$ by $[n]=\{1,\cdots,n\}$, randomly sample a subset of $n-r$ qubits $J=\{j_{1},\cdots,j_{n-r}\}$ with probability $p_{J}$. The set of left qubits is denoted by $K=\{k_{1},\cdots,k_{r}\}$.
@@
Figure 4: Illustration of the DPSO procedure for $n=4,\ell=1$. The left blue rectangle represents the $4$-qubit state $\rho$, while the right dashed rectangle represents the query model of the target state $|\psi\rangle$ in a classical computer. On the left side, the $2$nd qubit is sampled in the first step and Pauli-$Z$ measurements are applied to the remaining three qubits, yielding measurement outcomes $(z_{1},z_{3},z_{4})$. The classical shadow of the $2$nd qubit’s state, $\hat{\sigma}_{(z_{1},z_{3},z_{4})}$, is then constructed from the outcome of the randomized Pauli measurement. On the right side, given the measurement outcomes $(z_{1},z_{3},z_{4})$, the classical model is queried to obtain the amplitudes corresponding to $\tilde{\boldsymbol{z}}_{1}=(z_{1},0,z_{3},z_{4})$ and $\tilde{\boldsymbol{z}}_{2}=(z_{1},1,z_{3},z_{4})$. After constructing the projector $L_{(z_{1},z_{3},z_{4})}$, the shadow overlap is computed classically.
@@
where $|s_{a}\rangle$ is the eigenstate corresponding to the outcome $s_{a}$ of the Pauli measurement applied to qubit $k_{a}\in K$. By the definition of the classical shadow [20], there exists the equality
@@
with $|\widetilde{\phi}_{K,\boldsymbol{z}}\rangle:=\langle\boldsymbol{z}|\psi\rangle$. Hence, the shadow expectation of $\hat{\omega}$ is
@@
Figure 5: Illustration of the histograms of spectral gaps for the Haar random pure states with different qubits. The averaged spectral gaps are show in Fig. 3. Here, the first and second figures corresponds to the SOP and DPSO protocols respectively.
@@ IV.2 Sample complexity
$|\hat{\omega}|=$ $|\langle\hat{\zeta}_{K,\boldsymbol{z}},|\phi_{K,\boldsymbol{z}}\rangle\langle\phi_{K,\boldsymbol{z}}|\rangle_{\text{HS}}|$
@@
Compared with the sample complexity of the SOP in Eq.(24), the DPSO protocol’s sample complexity has the same dependence on parameters $\delta,\epsilon$ and the spectral gap. But if two levels are equal, $r=\ell$, the sample complexity of the DPSO protocol is smaller than that of the SOP by a factor $2^{2\ell-2}$. Further, just as illustrated in Fig. 3 and Fig. 5, the spectral gap $\nu(\Omega)$ may have larger value than $\nu(L)$ with high probability. When $r=\ell=1$, it’s still probable that the DPSO protocol has a smaller sample complexity than the SOP, since the averaged spectral gap of DPSO is much smaller than that of SOP as illustrated in Fig. 3. Given a specific target state $|\psi\rangle$, if the level $r$, $\delta$ and $\epsilon$ are fixed, the sample complexity only depends on the spectral gap of $\Omega$. Hence, to find the optimal strategy operator $\Omega$ in the DPSO protocol 3, we need to solve the following optimization problem
@@ V Stabilizer formalism
In order to illustrate the DPSO protocol, we apply it to the stabilizer state. We can show that our protocol can deal with the stabilizer state, including the GHZ state, more naturally than the SOP. The problem of verification of the stabilizer state has been solved in the PLM framework elegantly [2, 8], and the protocols are optimal or near-optimal. So we don’t seek to find an explicit optimal DPSO protocol for verifying a specific stabilizer state but provide a method to determine the optimal strategy operator.
@@ V.1 Pauli measurement and the test operator
Considering the DPSO protocol 3, let’s focus on a composite Pauli measurement $\mu\in\mathbb{Z}_{2}^{2n}$ with weight $t=n-r$ and assume the non-trivial single Pauli measurements are applied to the set of qubits $J=\{j_{1},\cdots,j_{t}\}$. There are $t$ non-trivial Pauli operators like $\mathrm{i}^{\mu_{j_{a}}^{\mathrm{x}}\mu_{j_{a}}^{\mathrm{z}}}X_{j_{a}}^{\mu_{j_{a}}^{\mathrm{x}}}Z_{j_{a}}^{\mu_{j_{a}}^{\mathrm{z}}}$ with $a=1,\cdots,t$, then these Pauli operators generate a stabilizer group $\mathcal{T}_{\mu}$ with cardinality $|\mathcal{T}_{\mu}|=2^{t}$. Applying the composite Pauli measurement to a state $\rho$, the outcome of qubit $j_{a}$ is $(-1)^{v_{a}}$ with $v_{a}\in\{0,1\}$, then the measurement outcomes of $t$ qubits correspond to a $t$-dimensional vector $\boldsymbol{v}\in\mathbb{Z}_{2}^{t}$. For each vector $\boldsymbol{v}\in\mathbb{Z}_{2}^{k}$, we can construct a new stabilizer groups $\mathcal{T}_{\mu,\boldsymbol{v}}$ generated by such Pauli operator $(-1)^{v_{a}}\mathrm{i}^{\mu_{j_{a}}^{\mathrm{x}}\mu_{j_{a}}^{\mathrm{z}}}X_{j_{a}}^{\mu_{j_{a}}^{\mathrm{x}}}Z_{j_{a}}^{\mu_{j_{a}}^{\mathrm{z}}}$ with $a=1,\cdots,t$, its corresponding projector is
@@
Note that the reduced states corresponding to some outcomes $\boldsymbol{v}$ may be 00, then such terms vanish and disappear in Eq.(43), so we don’t need to worry the denominator. The test operator $\Omega_{\mu}$ in Eq.(43) has another more useful form, as summarized in the following theorem.
@@
To prove Eq.(48), let’s first assume that $\rho$ is a stabilizer state with stabilizer group $\mathcal{S}$, then according to the results in [8], we know that
@@ V.2 The stabilizer state
The key point is that $\mathcal{R}_{\mu}$ is a stabilizer subgroup of $\mathcal{S}$, which has been known in [24, 25], then $\Omega_{\mu}$ is the projector into the stabilizer code of $\mathcal{R}_{\mu}$. This fact is summarized rigorously in the following corollary.
@@
From the above Corollary, we can easily derive the cardinality of the group, $|\mathcal{R}_{\mu}|=2^{n-t}|\mathcal{T}_{\mu}\cap\mathcal{S}|$. Compared to the similar results in [8], our results actually encompass theirs as a special case of $t=n$. For $\mathcal{R}_{\mu}$ is a subgroup of $\mathcal{S}$ and all elements of $\mathcal{S}$ commute with each other, therefore all test operators $\Omega_{\mu}$ commute with each other and can be diagonalizable simultaneously.
@@
with $\gamma_{\mu,\boldsymbol{w}}$ denoting the diagonal element. Further, the diagonal elements of $\Omega_{\mu}$ are either 00 or $1$,
@@
The above optimization problem can be further reduced into a linear programming as done in [8].
@@
Fortunately, the physical interesting target states have many symmetries, for example, the GHZ state. The symmetries of the target state can reduce the number of test operators very much. Furthermore, taking into the consideration of the symmetries of the target state can also boost the spectral gap of the strategy operator [6, 9].
Actually, there have been several explicit protocols for verifying the stabilizer state in [2, 5, 8]. So compared with finding the optimal strategy operator, sometimes we are more interested in the special strategy operator $\Omega_{\text{u}}$ corresponding to the uniform sampling, since we don’t need to solve an optimization problem. For example, without considering the symmetries of the target state, we let all $p_{\mu}$’s are equal, i.e.,
@@ V.3 The GHZ state
It’s obvious that $|\text{GHZ}_{n}\rangle$ is invariant under the action of any permutation $P\in\mathcal{P}_{n}$ of $n$ qubits, then if two different composite Pauli measurements are the same up to a permutation of $n$ qubits, they don’t result in any difference in the physical sense, hence the probabilities $p_{\mu}$ for sampling such two measurements should be the same. However, for the labeling of $n$ qubits, the test operators for the composite Pauli measurements being the same up to some permutations actually are different, so we need to take the average of them together. More explicitly, two symplectic vectors $\mu,\mu^{\prime}$ labeling two composite Pauli measurements are equivalent, denoted as $\mu\sim\mu^{\prime}$, if they can be transformed into each other by applying some permutation $P\in\mathcal{P}_{n}$. Each equivalence class $[\mu]$ can be distinguished by the number of Pauli-$X$, $Y$ and $Z$ measurements, $(m_{x},m_{y},m_{z})$. For each equivalence class $[\mu]$ labelled by $(m_{x},m_{y},m_{z})$, we sum over all corresponding test operators and divide it by the number of test operators in this class as
@@ V.3.1 $3$-qubit GHZ state
Level-$1$
According to the previous discussions, we need to compute all test operators $\Omega_{\mu}$ with $\mu$ having weight $2$. But for simplicity, we can only compute the expressions of $\Omega_{K,\boldsymbol{l}}$ with $K=\{3\}$ ($J=\{1,2\}$) and $\boldsymbol{l}\in\{1,2,3\}^{2}$, the other test operators can be obtained by permutations. Having obtained all test operators, we can consider two strategy operators corresponding to two different uniform smapling.
  • The first one is the strategy operator corresponding to the naively uniform sampling, the expressions of whose eigenvalues are given in Eq.(69). In this case, we compute the spectral gap is $4/9$.
  • The second one is the strategy operator corresponding to the uniform sampling over the equivalence classes. There are totally $6$ classes: $(m_{x},m_{y},m_{z})=(2,0,0)$, $(0,2,0)$, $(0,0,2)$, $(1,1,0)$, $(1,0,1)$ or $(0,1,1)$. Finally, we get the spectral gap is $1/2$.
Level-$2$
In the level-$2$ protocol, we need to compute the test operators $\Omega_{\mu}$ with $\mu$ having weight $1$. As before, we consider two cases of uniform sampling. For the naive uniform sampling, we can compute the spectral gap is $2/3$. While after considering the symmetries of $|\mathrm{GHZ}_{3}\rangle$, we obtain the spectral gap as $2/3$ too.

### v4

@@ IV Directly partial shadow overlap protocol
Figure 3: Illustration of the averaged spectral gaps of $\nu(L)$ and $\nu(\Omega)$ for the SOP and DPSO protocol. The target states are Haar-random pure states, and the averaged spectral gaps are computed over 10001000 samples. The first figure shows that the averaged spectral gaps decreases as the number of qubits increases, and that averaged $\nu(\Omega)$ is consistently larger than the averaged $\nu(L)$. The second figure illustrates the averaged spectral gaps increase with the protocol level, and the averaged spectral gap $\nu(\Omega)$ exceeds the averaged $\nu(L)$ across all levels.
We have discussed in detail the differences between the PLM framework and the SOP. In the SOP, only Pauli-Z measurements are applied to nrn-r qubits, which prevents the direct application of the SOP to certain states, such as GHZ states. Additionally, the operator $L$ appears somewhat unnatural and complex. To address these limitations, we aim to generalize the protocol by incorporating other types of Pauli measurements and selecting a more natural operator than $L$. Inspired by measurements used in Li and Zhu (2025), we propose a new protocol, called as the directly partial shadow overlap (DPSO) protocol, which follows the same core principles as the SOP and utilizes the classical shadow method.
@@ IV.1 Description of the verification procedure
In general, our DPSO protocol is similar with the SOP in Huang et al. (2024a) and both of them are based on the i.i.d. assumption for the states produced by the quantum device. As outlined in Protocol 3, we need to run $N$ times of Algorithm 2 to compute $N$ shadow overlaps $\hat{\omega}_{a}$. In Algorithm 2, the steps $1,2,3$ correspond to the first stage of the SOP, while the steps $4,5$ correspond to the second stage of classical post-processing. The procedure of Algorithm 2 is also illustrated in Figure 4.
@@
  • For the randomly selected nrn-r qubits, we randomly apply Pauli-$X$, $Y$, or $Z$ measurements, rather than only the Pauli-$Z$ measurements.
@@
The last two differences are more important. The second modification enables us to handle a wider range of states with special structures, such as the GHZ states and stabilizer states. The third adjustment opens the possibility of constructing the optimal protocol, where the strategy operator achieves the optimal spectral gap. Hence, when the level $r=1$ and only Pauli-$Z$ measurements are applied, the DPSO protocol reduces to the SOP and is similar with the protocol in Li and Zhu (2025), then the DPSO protocol can also verify almost all Haar random pure states, just as the SOP Huang et al. (2024a). Finally, it is important to emphasize that the measurement applied in each run of the DPSO protocol is not an LOCC measurement, but rather a randomized Pauli measurement. Therefore, the DPSO protocol should not be regarded as an adaptive version of the SOP.
@@
1:Label $n$ qubits of state $\rho$ by $n]=\{1,\cdots,n\}$, randomly sample a subset of nrn-r qubits $J=\{j_{1},\cdots,j_{n-r}\}$ with probability $p_{J}$. The set of remaining qubits is denoted by $K=\{k_{1},\cdots,k_{r}\}$.
@@
Figure 4: Illustration of the DPSO procedure for $n=4,\ell=1$. The left blue rectangle represents the $4$-qubit state $\rho$, while the right dashed rectangle represents the query model of the target state $|\psi\rangle$ in a classical computer. On the left side, the 22nd qubit is sampled in the first step and Pauli-$Z$ measurements are applied to the remaining three qubits, yielding measurement outcomes $(z_{1},z_{3},z_{4})$. The classical shadow of the 22nd qubit’s state, $\hat{\sigma}_{(z_{1},z_{3},z_{4})}$, is then constructed from the outcome of the randomized Pauli measurement. On the right side, given the measurement outcomes $(z_{1},z_{3},z_{4})$, the classical model is queried to obtain the amplitudes corresponding to $\tilde{\boldsymbol{z}}_{1}=(z_{1},0,z_{3},z_{4})$ and $\tilde{\boldsymbol{z}}_{2}=(z_{1},1,z_{3},z_{4})$. After constructing the projector $L_{(z_{1},z_{3},z_{4})}$, the shadow overlap is computed classically.
@@
where $|s_{a}\rangle$ is the eigenstate corresponding to the outcome $s_{a}$ of the Pauli measurement applied to qubit $k_{a}\in K$. By the definition of the classical shadow Huang et al. (2020), there exists the equality
@@
with $|\widetilde{\phi}_{K,\boldsymbol{z}}\rangle:=\langle\boldsymbol{z}|\psi\rangle$. The computation of $|\phi_{K,\boldsymbol{z}}\rangle$ can be efficient for certain classical descriptions of the target state, such as when $|\psi\rangle$ is represented as a matrix product state, or under a more powerful query model than that considered in Huang et al. (2020). Further discussion of this topic is provided in Appendix A. Hence, the shadow expectation of $\hat{\omega}$ is
@@
Figure 5: Illustration of the histograms of spectral gaps for the Haar random pure states with different number of qubits. The averaged spectral gaps are shown in Fig. 3. Here, the first and second figures correspond to the SOP and DPSO protocols respectively.
@@ IV.2 Sample complexity
[…]
@@
Compared with the sample complexity of the SOP in Eq.(24), the DPSO protocol’s sample complexity has the same dependence on parameters $\delta,\epsilon$ and the spectral gap. But if two levels are equal, $r=\ell$, the sample complexity of the DPSO protocol is smaller than that of the SOP by a factor $2^{2\ell-2}$. Further, just as illustrated in Fig. 3 and Fig. 5, the spectral gap $\nu(\Omega)$ may have larger value than $\nu(L)$ with high probability. When $r=\ell=1$, it’s still probable that the DPSO protocol has a smaller sample complexity than the SOP, since the averaged spectral gap of DPSO is much smaller than that of SOP as illustrated in Fig. 3. Moreover, the DPSO protocol can also be modified analogously to the modification of Protocol $1$ in Ref. Huang et al. (2024b), in order to achieve an improved dependence for the sample complexity as $\mathcal{O}(\epsilon^{-1}\nu(L)^{-1})$.
Given a specific target state $|\psi\rangle$, if the level $r$, $\delta$ and $\epsilon$ are fixed, the sample complexity only depends on the spectral gap of $\Omega$. Hence, to find the optimal strategy operator $\Omega$ in the DPSO protocol 3, we need to solve the following optimization problem
@@ V Stabilizer formalism
In order to illustrate the DPSO protocol, we apply it to the stabilizer state. We can show that our protocol can deal with the stabilizer state, including the GHZ state, more naturally than the SOP. The problem of verification of the stabilizer state has been solved in the PLM framework elegantly Pallister et al. (2018); Dangniam et al. (2020), and the protocols are optimal or near-optimal. So we don’t seek to find an explicit optimal DPSO protocol for verifying a specific stabilizer state but provide a method to determine the optimal strategy operator.
@@ V.1 Pauli measurement and the test operator
Considering the DPSO protocol 3, let’s focus on a composite Pauli measurement $\mu\in\mathbb{Z}_{2}^{2n}$ with weight t=nrt=n-r and assume the non-trivial single Pauli measurements are applied to the set of qubits $J=\{j_{1},\cdots,j_{t}\}$. There are $t$ non-trivial Pauli operators like $\mathrm{i}^{\mu_{j_{a}}^{\mathrm{x}}\mu_{j_{a}}^{\mathrm{z}}}X_{j_{a}}^{\mu_{j_{a}}^{\mathrm{x}}}Z_{j_{a}}^{\mu_{j_{a}}^{\mathrm{z}}}$ with $a=1,\cdots,t$, then these Pauli operators generate a stabilizer group $\mathcal{T}_{\mu}$ with cardinality $|\mathcal{T}_{\mu}|=2^{t}$. Applying the composite Pauli measurement to a state $\rho$, the outcome of qubit $j_{a}$ is $(-1)^{v_{a}}$ with $v_{a}\in\{0,1\}$, then the measurement outcomes of $t$ qubits correspond to a $t$-dimensional vector $\boldsymbol{v}\in\mathbb{Z}_{2}^{t}$. For each vector $\boldsymbol{v}\in\mathbb{Z}_{2}^{k}$, we can construct a new stabilizer groups $\mathcal{T}_{\mu,\boldsymbol{v}}$ generated by such Pauli operator $(-1)^{v_{a}}\mathrm{i}^{\mu_{j_{a}}^{\mathrm{x}}\mu_{j_{a}}^{\mathrm{z}}}X_{j_{a}}^{\mu_{j_{a}}^{\mathrm{x}}}Z_{j_{a}}^{\mu_{j_{a}}^{\mathrm{z}}}$ with $a=1,\cdots,t$, its corresponding projector is
@@
Note that the reduced states corresponding to some outcomes $\boldsymbol{v}$ may be 0, then such terms vanish and disappear in Eq.(43), so we don’t need to worry the denominator. The test operator $\Omega_{\mu}$ in Eq.(43) has another more useful form, as summarized in the following theorem.
@@
To prove Eq.(48), let’s first assume that $\rho$ is a stabilizer state with stabilizer group $\mathcal{S}$, then according to the results in Dangniam et al. (2020), we know that
@@ V.2 The stabilizer state
The key point is that $\mathcal{R}_{\mu}$ is a stabilizer subgroup of $\mathcal{S}$, which has been known in Hayden et al. (2016); Gross et al. (2021), then $\Omega_{\mu}$ is the projector into the stabilizer code of $\mathcal{R}_{\mu}$. This fact is summarized rigorously in the following corollary.
@@
From the above Corollary, we can easily derive the cardinality of the group, $|\mathcal{R}_{\mu}|=2^{n-t}|\mathcal{T}_{\mu}\cap\mathcal{S}|$. Compared to the similar results in Dangniam et al. (2020), our results actually encompass theirs as a special case of $t=n$. For $\mathcal{R}_{\mu}$ is a subgroup of $\mathcal{S}$ and all elements of $\mathcal{S}$ commute with each other, therefore all test operators $\Omega_{\mu}$ commute with each other and can be diagonalizable simultaneously.
@@
with $\gamma_{\mu,\boldsymbol{w}}$ denoting the diagonal element. Further, the diagonal elements of $\Omega_{\mu}$ are either 0 or $1$,
@@
The above optimization problem can be further reduced into a linear programming as done in Dangniam et al. (2020).
@@
Fortunately, the physical interesting target states have many symmetries, for example, the GHZ state. The symmetries of the target state can reduce the number of test operators significantly. Furthermore, taking into the consideration of the symmetries of the target state can also boost the spectral gap of the strategy operator Yu et al. (2019); Liu et al. (2019).
Actually, there have been several explicit protocols for verifying the stabilizer state in Pallister et al. (2018); Zhu and Hayashi (2019a); Dangniam et al. (2020). So compared with finding the optimal strategy operator, sometimes we are more interested in the special strategy operator $\Omega_{\text{u}}$ corresponding to the uniform sampling, since we don’t need to solve an optimization problem. For example, without considering the symmetries of the target state, we let all $p_{\mu}$’s are equal, i.e.,
@@ V.3 The GHZ state
It’s obvious that $|\text{GHZ}_{n}\rangle$ is invariant under the action of any permutation $P\in\mathcal{P}_{n}$ of $n$ qubits, then if two different composite Pauli measurements are the same up to a permutation of $n$ qubits, they don’t result in any difference in the physical sense, hence the probabilities $p_{\mu}$ for sampling such two measurements should be the same. However, for the labeling of $n$ qubits, the test operators for the composite Pauli measurements being the same up to some permutations actually are different, so we need to take the average of them together. More explicitly, two symplectic vectors $\mu,\mu^{\prime}$ labeling two composite Pauli measurements are equivalent, denoted as $\mu\sim\mu^{\prime}$, if they can be transformed into each other by applying some permutation $P\in\mathcal{P}_{n}$. Each equivalence class $\mu]$ can be distinguished by the number of Pauli-$X$, $Y$ and $Z$ measurements, $(m_{x},m_{y},m_{z})$. For each equivalence class $\mu]$ labelled by $(m_{x},m_{y},m_{z})$, we sum over all corresponding test operators and divide it by the number of test operators in this class as
@@ V.3.1 $3$-qubit GHZ state
Level-11
According to the previous discussions, we need to compute all test operators $\Omega_{\mu}$ with $\mu$ having weight $2$. For simplicity, we only compute the expressions of $\Omega_{K,\boldsymbol{l}}$ with $K=\{3\}$ ($J=\{1,2\}$) and $\boldsymbol{l}\in\{1,2,3\}^{2}$. The remaining test operators can be obtained through symmetry considerations via appropriate index permutations. Having obtained all test operators, we can consider two strategy operators corresponding to two different uniform sampling.
  • The first one is the strategy operator corresponding to the naively uniform sampling, the expressions of whose eigenvalues are given in Eq.(69). In this case, we conclude the spectral gap is $4/9$.
  • The second one is the strategy operator corresponding to the uniform sampling over the equivalence classes. There are totally $6$ classes: $(m_{x},m_{y},m_{z})=(2,0,0)$, $(0,2,0)$, $(0,0,2)$, $(1,1,0)$, $(1,0,1)$ or $(0,1,1)$. Finally, we conclude the spectral gap is $1/2$.
Level-22
In the level-$2$ protocol, we need to compute the test operators $\Omega_{\mu}$ with $\mu$ having weight $1$. As before, we consider two cases of uniform sampling. For the naive uniform sampling, we can conclude the spectral gap is $2/3$. While after considering the symmetries of $|\mathrm{GHZ}_{3}\rangle$, we obtain the spectral gap as $2/3$ too.